OpenStudy (anonymous):

Which decimal is equivalent to 8 9/12 ?

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

I give medals.

OpenStudy (danjs):

reduce 9/12 down

OpenStudy (danjs):

divide both top and bottom by 3

OpenStudy (anonymous):

K

OpenStudy (utterly_confuzzled):

So, take 9 and divide it by 12. the answer will be 8. and then whatever the answer of 9/12 is

OpenStudy (danjs):

what do you get after you reduce the fraction kdenice

TheSmartOne (thesmartone):

And we need choices...

OpenStudy (anonymous):

8.34?

OpenStudy (danjs):

9/12 = 3/4

TheSmartOne (thesmartone):

If there are choices..

OpenStudy (danjs):

3/4 is an easy decimal to remember 0.75

OpenStudy (anonymous):

A. 8.12 B. 8.34 C. 8.75 D. 8.9

OpenStudy (danjs):

8 9/12 = 8 3/4 = 8.75

OpenStudy (utterly_confuzzled):

3/4 = .75 Therefore the answer will be C 8.75

OpenStudy (danjs):

or if you want 3/4 = 75/100 = 0.75

OpenStudy (danjs):

right

OpenStudy (anonymous):

Ok!!!

OpenStudy (danjs):

are you also doing the way where you have to multiply numbers by say 100 or 1000, to turn fractions to decimals?

OpenStudy (danjs):

here is a little thing i just found, may help you with these http://www.mathsisfun.com/converting-fractions-decimals.html

OpenStudy (anonymous):

Thanks

OpenStudy (danjs):

that basicly has you put the denominator to multiples of 10, like 100 1000 10000, so it is easy to see what the decimal is